Find 4 consecutive even integers such that four less than twice the samllest is equal to three times the largest, increased by two.
Answer by ankor@dixie-net.com(15652)

Looking over this problem it, becomes apparent that the numbers are negative
:
Find 4 consecutive even integers
x, (x-2), (x-4), (x-6); where x is the greatest number, (x-6) is the smallest
;
such that four less than twice the smallest is equal to three times the largest, increased by two.
:
2(x-6) - 4 = 3x + 2
:
2x - 12 - 4 = 3x + 2
:
2x - 16 = 3x + 2
:
2x - 3x = 2 + 16
:
-x = +18
:
x = -18
:
The numbers, smallest to greatest: -24, -22, -20, -18
:
:
Check solution using the statement:
"four less than twice the smallest is equal to three times the largest, increased by two.:
2(-24) - 4 = 3(-18) + 2
-48 - 4 = -54 + 2; confirms our solution